imqwerty (imqwerty):

100ml H2SO4 solution having molarity 1M and density 1.5g/ml is mixed with 400ml of water. Calculate final molarity of H2SO4 solution, if final density is 1.25g/ml.

OpenStudy (hwyl):

first, make sure you have a good grasp of the definition of molarity

OpenStudy (adi3):

@imqwerty

imqwerty (imqwerty):

well i know how to solve the question but i think the question is wrong

OpenStudy (adi3):

well then type the correct question

OpenStudy (hwyl):

then show your solution

imqwerty (imqwerty):

ok 1st i'll show the solution and then i'll show how the question is wrong

imqwerty (imqwerty):

Molarity=moles/volume(in lts.) for h2so4 :- 1M=n(moles)/0.1 n=0.1 total weight of final solution = 100 x 1.5 +400 = 550gm volume of final solution = 550/1.25 = 440ml = 0.440L molarity of h2so4 in final solution = 0.1/0.440 = 0.227M

imqwerty (imqwerty):

it is given that the final density is 1.25g/ml the final mix has h2so4 and h2o total mass of final mixture as calculated above is = 550gm total volume of final mix. = vol. of h2so4 + vol. of h2o = 100+400 = 500 so final density of final solution = mass/volume = 550/500 = 1.1gm/ml but in the question the final density given is 1.25gm/ml now how is that possible??

OpenStudy (hwyl):

when you dilute a solution, does the density increase or decrease?

imqwerty (imqwerty):

it increases??

OpenStudy (hwyl):

prove it mathematically

imqwerty (imqwerty):

well i don't know how to do so??

OpenStudy (hwyl):

can you use the density formula?

imqwerty (imqwerty):

yes

OpenStudy (hwyl):

then try

imqwerty (imqwerty):

the way i showed that the final density is 1.1 is also correct.

imqwerty (imqwerty):

@hwyl how will you prove that density is not 1.1 and its 1.25??

OpenStudy (anonymous):

You are meant to calculate the concentration of H2SO4 and then to use that to find the final density of the solution and CHECK that it is 1.25.

OpenStudy (anonymous):

Conentration = molarity

imqwerty (imqwerty):

@ShizukaTheOtaku no we don't have to check whether the density given is correct or not rather we are given the density to find the molarity. But i wished to check whether the density given was correct or not and in the calculation above which i posted the density is comming out to be different from what is given in the question.

imqwerty:

Little did I know

imqwerty:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @imqwerty it is given that the final density is 1.25g/ml the final mix has h2so4 and h2o total mass of final mixture as calculated above is = 550gm total volume of final mix. = vol. of h2so4 + vol. of h2o = 100+400 = 500 so final density of final solution = mass/volume = 550/500 = 1.1gm/ml but in the question the final density given is 1.25gm/ml now how is that possible?? \(\color{#0cbb34}{\text{End of Quote}}\) Good observation, but the final volume that you've taken is wrong. It should be calculated by the mass/density formula and not by just adding volH2O and volH2SO4, because the volumes can change after the reaction but the mass remains constant.

imqwerty:

:)

imqwerty:

@imqwerty wrote:
it increases??
bruh...

imqwerty:

@imqwerty wrote:
\(\color{#0cbb34}{\text{Originally Posted by}}\) @imqwerty it is given that the final density is 1.25g/ml the final mix has h2so4 and h2o total mass of final mixture as calculated above is = 550gm total volume of final mix. = vol. of h2so4 + vol. of h2o = 100+400 = 500 so final density of final solution = mass/volume = 550/500 = 1.1gm/ml but in the question the final density given is 1.25gm/ml now how is that possible?? \(\color{#0cbb34}{\text{End of Quote}}\) Good observation, but the final volume that you've taken is wrong. It should be calculated by the mass/density formula and not by just adding volH2O and volH2SO4, because the volumes can change after the reaction but the mass remains constant.
cool, but doesn't answer "why" exactly is the volume lesser than expected.